Grade 8General Physics

in a factory it is desired to lift 2000 kg of metal through a distance of 12 m in 1 minute. find the minimum horse power of the engine to be used.

prasanjeet kumar
12 Years agoGrade 8

1 Answer

reeease
9 Years ago
m=2000 kg
s=12 m
t=60 sec
p= w/t
p= ma/t      (w=ma)
p=mgh/t      (a=gh)
p=2000*9.8*12/60
p=3920 watt.
horse power = 3920/746
                   = 5.3 hp